A hurricane strikes and a rural area is without food or water. Three crews arrive. One can dispense needed supplies in 13 hours, a second in 16 hours, and a third in 22 hours. How long will it take all three crews working together to dispense food and water?
Answer by edjones(8007)

1st crew does 1/13 of the work in 1 hr.
2nd crew does 1/16 of the work in 1 hr.
3rd crew does 1/22 of the work in 1 hr.
Together 1/13 + 1/16 + 1/22 = 176/2288 + 143/2288 +104/2288 = 423/2288 of the work in 1 hr. (2288 is the LCD)
2288/423=5.4... hrs it takes the 3 crews to do the job.
.
Ed